# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ C.2 Setting and Notation
Under Assumption C.8, Theorem 6.7 implies that w.p. at least $1-\lambda$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$p_{\mathcal{S}}\geq\exp{\left(d_{1}^{\star}d_{0}\log\left(\sin\left(\alpha\right)\right)+d_{1}\log\left(\sin\left(\gamma\right)\right)+O\left(d_{1}^{\star}\log{\left(d_{0}\right)}+\log{\left(d_{1}\right)}\right)\right)}\,.$
We denote this event by $\mathcal{E}_{1}$. Recalling that $\hat{C}^{\mathrm{cont}}\geq-\log\left(p_{\mathcal{S}}\right)$ when conditioned on $\mathcal{E}_{1}$, from Prop. A.19 we deduce that,
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\,.$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)$ $=\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\right)+\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{2}\right)-\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\cup\mathcal{E}_{2}\right)$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\ -\lambda\,.$
Appendix D Proofs for Sparsest Quantized Interpolator Learning Rule
D.1 Setting and Notation
@@
Definition D.1.
$\mathcal{A}_{0}\left(\mathcal{S}\right)=h_{\bm{\theta}_{0}}$ returns the sparsest quantized interpolator,
$\bm{\theta}_{0}=\operatorname*{\arg\!\min}_{\bm{\theta}\in\mathcal{Q}^{M}}{\left\|\bm{\theta}\right\|_{0}}\;\text{s.t}\;\forall n\in\left[N\right]\;y^{\left(n\right)}h_{\bm{\theta}}\left(\mathbf{x}^{\left(n\right)}\right)>0\,,$
where $\left\|\bm{\theta}\right\|_{0}$ is the number of nonzero values in $\bm{\theta}$. With some abuse of notation, we use $\left\|\mathcal{A}_{0}\left(\mathcal{S}\right)\right\|_{0}$ and $\left\|\bm{\theta}_{0}\right\|_{0}$ interchangeably.
Recall that we denote the total number of parameters in a teacher network $h_{\bm{\theta}^{\star}}$ by
@@
the maximal number of non-zero weights and biases in $h_{\bm{\theta}^{\star}}$, respectively. Denote the class of fully-connected neural networks with at most $W\left(D^{\star}\right)$ non-zero weights and $B\left(D^{\star}\right)$ biases as $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$  [^11] This is different from $\mathcal{H}_{D}^{FC}$ as no specific depth and hidden layer widths are assumed for $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$. Notice that the number of neural networks in $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$ is bounded by the number of neural networks with $M\left(D^{\star}\right)$ edges and no bias terms. We denote the set of such neural networks by $\mathcal{H}_{M\left(D^{\star}\right)}$, then
@@
D.2 Generalization Bound
Lemma D.2.
@@ D.2 Generalization Bound
Since $h_{\bm{\theta}^{\star}}$ is an interpolating solution, we have that $\left\|\mathcal{A}_{0}\left(\mathcal{S}\right)\right\|_{0}\leq M\left(D^{\star}\right)$. The bound follows from the case in which all neurons have 0 bias, and the number of non-zero weights is equal to the number of non-zero parameters. ∎
@@
Collecting the bounds from (61), (62) and Lemma D.4 we find the following corollaries.
Corollary D.5.
@@
Using Theorem A.1 we get,
Corollary D.6.
Let $\epsilon>0$ and $\delta\in\left(0,1\right)$. With probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{0}\left(\mathcal{S}\right))\leq\epsilon$
@@
$N\geq\frac{2M\left(D^{\star}\right)\log\left(M\left(D^{\star}\right)+d_{0}\right)+M\left(D^{\star}\right)\log\left(Q\right)+\log\left(\frac{1}{\delta}\right)}{\epsilon}\,.$

### v2

@@ D.2 Setting and Notation
Under Assumption D.8, Theorem 5.6 implies that w.p. at least $1-\lambda$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\hat{p}_{\mathcal{S}}\geq\exp{\left(d_{1}^{\star}d_{0}\log\left(\sin\left(\alpha\right)\right)+d_{1}\log\left(\sin\left(\gamma\right)\right)+O\left(d_{1}^{\star}\log{\left(d_{0}\right)}+\log{\left(d_{1}\right)}\right)\right)}\,.$
We denote this event by $\mathcal{E}_{1}$. Recalling that $\hat{C}^{\mathrm{cont}}\geq-\log\left(\hat{p}_{\mathcal{S}}\right)$ when conditioned on $\mathcal{E}_{1}$, from Prop. B.23 we deduce that,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\,.$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)$ $=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\right)+\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{2}\right)-\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{E}_{1}\cup\mathcal{E}_{2}\right)$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\ -\lambda\,.$
Appendix E Proofs for Sparsest Quantized Interpolator Learning Rule
E.1 Setting and Notation
@@
Definition E.1.
$\mathcal{A}_{0}\left(\mathcal{S}\right)=h_{\boldsymbol{\theta}_{0}}$ returns the sparsest quantized interpolator,
$\boldsymbol{\theta}_{0}=\operatorname*{\arg\!\min}_{\boldsymbol{\theta}\in\mathcal{Q}^{M}}{\left\|\boldsymbol{\theta}\right\|_{0}}\;\text{s.t}\;\forall n\in\left[N\right]\;y^{\left(n\right)}h_{\boldsymbol{\theta}}\left(\mathbf{x}^{\left(n\right)}\right)>0\,,$
where $\left\|\boldsymbol{\theta}\right\|_{0}$ is the number of nonzero values in $\boldsymbol{\theta}$. With some abuse of notation, we use $\left\|\mathcal{A}_{0}\left(\mathcal{S}\right)\right\|_{0}$ and $\left\|\boldsymbol{\theta}_{0}\right\|_{0}$ interchangeably.
Recall that we denote the total number of parameters in a teacher network $h_{\boldsymbol{\theta}^{\star}}$ by
@@
the maximal number of non-zero weights and biases in $h_{\boldsymbol{\theta}^{\star}}$, respectively. Denote the class of fully-connected neural networks with at most $W\left(D^{\star}\right)$ non-zero weights and $B\left(D^{\star}\right)$ biases as $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$  [^13] This is different from $\mathcal{H}_{D}^{FC}$ as no specific depth and hidden layer widths are assumed for $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$. Notice that the number of neural networks in $\mathcal{H}_{W\left(D^{\star}\right),B\left(D^{\star}\right)}$ is bounded by the number of neural networks with $M\left(D^{\star}\right)$ edges and no bias terms. We denote the set of such neural networks by $\mathcal{H}_{M\left(D^{\star}\right)}$, then
@@
E.2 Generalization Bound
Lemma E.2.
@@ E.2 Generalization Bound
Since $h_{\boldsymbol{\theta}^{\star}}$ is an interpolating solution, we have that $\left\|\mathcal{A}_{0}\left(\mathcal{S}\right)\right\|_{0}\leq M\left(D^{\star}\right)$. The bound follows from the case in which all neurons have 0 bias, and the number of non-zero weights is equal to the number of non-zero parameters. ∎
@@
Collecting the bounds from (62), (63) and Lemma E.4 we find the following corollaries.
Corollary E.5.
@@
Using Theorem B.1 we get,
Corollary E.6.
Let $\varepsilon>0$ and $\delta\in\left(0,1\right)$. With probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{0}\left(\mathcal{S}\right))\leq\varepsilon$
@@
$N\geq\frac{2M\left(D^{\star}\right)\log\left(M\left(D^{\star}\right)+d_{0}\right)+M\left(D^{\star}\right)\log\left(Q\right)+\log\left(\frac{1}{\delta}\right)}{\varepsilon}\,.$